solve the following expression for z. - 3/4 - 2/3 = 1/3
Answer by stanbon(75887)

solve the following expression for z.
- 3/4z - 2/3 = 1/3
-(3/4)z = 1
Multiply both sides by -4/3 to get:
z = -4/3
==============
Cheers,
Stan H.